OpenStudy (anonymous):

A parallelogram has an area of 42 cm2. What would the area be if the base was one-third as long and the height was twice as long?

OpenStudy (anonymous):

i figured it out... its 28

OpenStudy (anonymous):

yes you are correct this is the logic: A=bxh if b is shrinked to 1/3 and h is twice as long then you have A=(b/3)x(2h)=(2/3)bxh So your area is reduced to 2/3 of the area that the parallelogram with base b & height h